# Directional testing for one-way MANOVA in divergent dimensions

### v2

@@ B.3 Proof of Proposition 1
Figure S3: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n^{\tau}$ and $n=\sum_{i=1}^{3}100=300$, at nominal level $\alpha=0.05$ given by the gray horizontal line.
Table S4: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n^{\tau}$ and $n=\sum_{i=1}^{3}500=1500$, at nominal level $\alpha=0.05$.
@@
Figure S4: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n^{\tau}$ and $n=\sum_{i=1}^{3}500=1500$, at nominal level $\alpha=0.05$ given by the gray horizontal line.
S1.4 Robustness to misspecification
@@
More in detail, a multivariate $t$ distribution with location $0_{p}$, scale matrix $I_{p}$ and degrees of freedom $5$, a multivariate skew-normal distribution with location $1_{p}$, scale matrix $\Omega=(\omega_{jl})=(0.2)^{|j-l|}$ and shape parameter $1_{p}$, and a multivariate Laplace distribution with mean vector $1_{p}$ and identity covariance matrix.
For hypothesis (8) in the paper, Figures S5S6 and Tables S5S6 show the empirical size at the nominal level $\alpha=0.05$ if the underling distribution is misspecified. We see that the directional test still maintains the hightest accuracy.
Figure S5: Empirical size for the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper, at nominal level $\alpha=0.05$ given by the gray horizontal line. The left, middle and right panels correspond to multivariate $t$, multivariate skew-normal, and multivariate Laplace distributions of the true generating process, respectively, with $n_{i}=100$ and $g=3$.
Figure S6: Empirical size for the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two Skovgaard’s modifications (Sko1 and Sko2, respectively) for hypothesis (8) in the paper, at nominal level $\alpha=0.05$ given by the gray horizontal line. The left, middle and right panels correspond to multivariate $t$, multivariate skew-normal, and multivariate Laplace distributions of the true generating process, respectively, with $n_{i}=500$ and $g=3$.
Table S5: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=100$ and $g=3$, at nominal level $\alpha=0.05$.
@@
Table S6: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=500$ and $g=3$, at nominal level $\alpha=0.05$.
@@
Appendix S2 Simulation studies for Heteroscedastic one-way MANOVA
This section is studied the performance of directional test for heteroscedastic one-way MANOVA, comparing with LRT, Sko1 and Sko1. In particular, when the number of groups $g=2$, we also consider the $F$-approximation for the Behrens-Fisher test $T^{*2}$. The simulation results are computed via Monte Carlo simulation based on 10,000 replications.
S2.1 Empirical results for the moderate setup
Groups of size $n_{i},i\in\{1,\dots,g\}$, are generated from a $p$-variate standard normal distribution $N_{p}(0_{p},\Lambda_{i}^{-1})$ under the null hypothesis. We use an autoregressive structure for the covariance matrices. i.e. $\Lambda_{i}^{-1}=(\sigma_{jl})_{p\times p}=(\rho_{i}^{|j-l|})_{p\times p}$, with the $\rho_{i}$ chosen to an equally-distance sequence from 0.1 to 0.9 of length $g$. For each simulation experiment, we show results for $p=\lceil n_{i}^{\tau}\rceil$ with $\tau=j/24$, $j\in\{6,7,\cdots,22\}$ and $n_{i}\in\{100,500,1000\}$ and $k=2$.
Table S7: Empirical size of the directional test (DT), Behrens-Fisher test (BF) (Nel and Merwe, 1986), likelihood ratio test (LRT), and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (9) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=100,500,1000$, at nominal level $\alpha=0.05$.
@@
S2.2 Robustness to misspecification
Table S8: Empirical size of the directional test (DT), Behrens-Fisher test (BF) (Nel and Merwe, 1986), likelihood ratio test (LRT), and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (9) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=100$ and $g=2$, at nominal level $\alpha=0.05$.
@@
Table S9: Empirical size of the directional test (DT), Behrens-Fisher test (BF) (Nel and Merwe, 1986), likelihood ratio test (LRT), and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (9) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=500$ and $g=2$, at nominal level $\alpha=0.05$.
@@
S2.3 Empirical results for large number $k$ of groups
Table S10: Empirical size of the directional test (DT), likelihood ratio test (LRT), and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (9) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=100$ and $g=30$, at nominal level $\alpha=0.05$.
$\tau$ ($p$) DT LRT Sko1 Sko2
0.250 (4) 0.050 0.081 0.050 0.050
0.333 (5) 0.050 0.091 0.050 0.050
0.417 (7) 0.049 0.122 0.050 0.049
0.500 (10) 0.050 0.186 0.051 0.050
0.583 (15) 0.054 0.391 0.058 0.054
0.667 (22) 0.047 0.751 0.056 0.048
0.750 (32) 0.052 0.990 0.087 0.062
0.833 (47) 0.049 1.000 0.223 0.095
0.917 (69) 0.048 1.000 0.932 0.405
@@
Figure S7: Empirical size of the directional test (DT), likelihood ratio test (LRT), and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (9) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=100$ and $g=30$, at nominal level $\alpha=0.05$ given by the gray horizontal line.
Table S11: Empirical size of the directional test (DT), likelihood ratio test (LRT), and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (9) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=1000$ and $g=5$, at nominal level $\alpha=0.05$
$\tau$ ($p$) DT LRT Sko1 Sko2
0.250 (6) 0.053 0.055 0.053 0.053
0.333 (10) 0.055 0.058 0.055 0.055
0.417 (18) 0.048 0.054 0.048 0.048
0.500 (32) 0.048 0.061 0.048 0.048
0.583 (57) 0.051 0.089 0.052 0.052
0.667 (100) 0.054 0.160 0.055 0.054
0.750 (178) 0.053 0.462 0.058 0.055
0.833 (317) 0.050 0.983 0.083 0.063
@@
Figure S8: Empirical size of the directional test (DT), likelihood ratio test (LRT), and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (9) in the paper, with $p=n_{i}^{\tau}$ and $n_{i}=1000$ and $g=5$, at nominal level $\alpha=0.05$ given by the gray horizontal line.

### v3

@@ 6 Conclusions
Figure S4: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n^{\tau}$ and $n=\sum_{i=1}^{3}100=300$, at nominal level $\alpha=0.05$ given by the gray horizontal line.
Table S4: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n^{\tau}$ and $n=\sum_{i=1}^{3}500=1500$, at nominal level $\alpha=0.05$.
@@
Figure S5: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n^{\tau}$ and $n=\sum_{i=1}^{3}500=1500$, at nominal level $\alpha=0.05$ given by the gray horizontal line.
S3.5 Robustness to misspecification for high dimensional setup
In general, all the approaches examined so far rely on the assumed normal model and are not guaranteed to be robust under model misspecification. We can assess numerically the robustness of the various competitors using simulations. We consider three different distributions for the true generating process: multivariate $t$, multivariate skew-normal (azzalini1999statistical) or multivariate Laplace. In more detail, a multivariate $t$ distribution with location $0_{p}$, scale matrix $I_{p}$ and degrees of freedom $5$, a multivariate skew-normal distribution with location $1_{p}$, scale matrix $\Omega=(\omega_{jl})=(0.2)^{|j-l|}$ and shape parameter $1_{p}$, and a multivariate Laplace distribution with mean vector $1_{p}$ and identity covariance matrix. Simulation results are based on $10,000$ replications.
For the homoscedastic case, Figures S6S7 show the empirical size at the nominal level $\alpha=0.05$ if the distribution is misspecified. We see that the directional test still maintains the highest accuracy.
Figure S6: Empirical size for the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (11) with $g=3$, at nominal level $\alpha=0.05$ given by the dashed gray horizontal line. The left, middle and right panels correspond to multivariate $t$, multivariate skew-normal, and multivariate Laplace distributions of the true generating process, respectively, with $n_{i}=100$.
Figure S7: Empirical size for the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2, respectively) for hypothesis (11) with $g=3$, at nominal level $\alpha=0.05$ given by the dashed gray horizontal line. The left, middle and right panels correspond to multivariate $t$, multivariate skew-normal, and multivariate Laplace distributions of the true generating process, respectively, with $n_{i}=500$.
S3.6 Robustness to misspecification for moderate setup
@@
In more detail, a multivariate $t$ distribution with location $0_{p}$, scale matrix $I_{p}$ and degrees of freedom $5$, a multivariate skew-normal distribution with location $1_{p}$, scale matrix $\Omega=(\omega_{jl})=(0.2)^{|j-l|}$ and shape parameter $1_{p}$, and a multivariate Laplace distribution with mean vector $1_{p}$ and identity covariance matrix.
For hypothesis (8) in the paper, Figures S8S9 and Tables S5S6 show the empirical size at the nominal level $\alpha=0.05$ if the underling distribution is misspecified. We see that the directional test still maintains the hightest accuracy.
Figure S8: Empirical size for the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper, at nominal level $\alpha=0.05$ given by the gray horizontal line. The left, middle and right panels correspond to multivariate $t$, multivariate skew-normal, and multivariate Laplace distributions of the true generating process, respectively, with $n_{i}=100$ and $g=3$.
Figure S9: Empirical size for the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2, respectively) for hypothesis (8) in the paper, at nominal level $\alpha=0.05$ given by the gray horizontal line. The left, middle and right panels correspond to multivariate $t$, multivariate skew-normal, and multivariate Laplace distributions of the true generating process, respectively, with $n_{i}=500$ and $g=3$.
Table S5: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=100$ and $g=3$, at nominal level $\alpha=0.05$.
@@
Table S6: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=500$ and $g=3$, at nominal level $\alpha=0.05$.
@@
S4 Simulation studies for Heteroscedastic one-way MANOVA
This section is studied the performance of directional test for heteroscedastic one-way MANOVA, comparing with LRT, Sko1 and Sko2. In particular, when the number of groups $g=2$, we also consider the $F$-approximation for the Behrens-Fisher test $T^{*2}$. The simulation results are computed via Monte Carlo simulation based on 10,000 replications.
S4.1 Empirical results for the moderate setup
Groups of size $n_{i},i\in\{1,\dots,g\}$, are generated from a $p$-variate standard normal distribution $N_{p}(0_{p},\Lambda_{i}^{-1})$ under the null hypothesis. We use an autoregressive structure for the covariance matrices. i.e. $\Lambda_{i}^{-1}=(\sigma_{jl})_{p\times p}=(\rho_{i}^{|j-l|})_{p\times p}$, with the $\rho_{i}$ chosen to an equally-distance sequence from 0.1 to 0.9 of length $g$. For each simulation experiment, we show results for $p=\lceil n_{i}^{\tau}\rceil$ with $\tau=j/24$, $j\in\{6,7,\cdots,22\}$ and $n_{i}\in\{100,500,1000\}$ and $k=2$. Note that for $n_{i}=1000$ the simulations results are based on $5000$ replications when $j\in\{21,22\}$ due to the expensive computational cost.
Table S7: Empirical size of the directional test (DT), Behrens-Fisher test (BF) (nel1986), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (9) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=100,500,1000$, at nominal level $\alpha=0.05$.
@@
Figure S10: Empirical size of the directional test (DT), Behrens-Fisher test (BF) (nel1986), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (12) with $g=2$, at nominal level $\alpha=0.05$ given by the dashed gray horizontal line. The left, middle and right panels correspond to $n_{i}=1000$, respectively.
S4.2 Empirical results for large number $g$ of groups
Table S8: Empirical size of the directional test (DT), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (9) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=100$ and $g=30$, at nominal level $\alpha=0.05$.
$\tau$ ($p$) DT LRT Sko1 Sko2
0.250 (4) 0.050 0.081 0.050 0.050
0.333 (5) 0.050 0.091 0.050 0.050
0.417 (7) 0.049 0.122 0.050 0.049
0.500 (10) 0.050 0.186 0.051 0.050
0.583 (15) 0.054 0.391 0.058 0.054
0.667 (22) 0.047 0.751 0.056 0.048
0.750 (32) 0.052 0.990 0.087 0.062
0.833 (47) 0.049 1.000 0.223 0.095
0.917 (69) 0.048 1.000 0.932 0.405
Figure S11: Empirical size of the directional test (DT), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (9) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=100$ and $g=30$, at nominal level $\alpha=0.05$ given by the gray horizontal line.
Table S9: Empirical size of the directional test (DT), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (9) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=1000$ and $g=5$, at nominal level $\alpha=0.05$
$\tau$ ($p$) DT LRT Sko1 Sko2
0.250 (6) 0.053 0.055 0.053 0.053
0.333 (10) 0.055 0.058 0.055 0.055
0.417 (18) 0.048 0.054 0.048 0.048
0.500 (32) 0.048 0.061 0.048 0.048
0.583 (57) 0.051 0.089 0.052 0.052
0.667 (100) 0.054 0.160 0.055 0.054
0.750 (178) 0.053 0.462 0.058 0.055
0.833 (317) 0.050 0.983 0.083 0.063
Figure S12: Empirical size of the directional test (DT), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (9) in the paper, with $p=n_{i}^{\tau}$ and $n_{i}=1000$ and $g=5$, at nominal level $\alpha=0.05$ given by the gray horizontal line.
S4.3 Robustness to misspecification
Table S10: Empirical size of the directional test (DT), Behrens-Fisher test (BF) (nel1986), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (9) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=100$ and $g=2$, at nominal level $\alpha=0.05$.
@@
Table S11: Empirical size of the directional test (DT), Behrens-Fisher test (BF) (nel1986), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (9) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=500$ and $g=2$, at nominal level $\alpha=0.05$.
@@
Figure S13: Empirical size for the directional test (DT), Behrens-Fisher test (BF), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (9) with $g=2$, at nominal level $\alpha=0.05$ given by the dashed gray horizontal line. The left, middle and right panels correspond to multivariate $t$, multivariate skew-normal, and multivariate Laplace distributions if the true generating process, respectively, with $n_{i}=100$.
@@
Figure S14: Empirical size for the directional test (DT), Behrens-Fisher test (BF), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (9), at nominal level $\alpha=0.05$ given by the gray horizontal line. The left, middle and right panels correspond to multivariate $t$, multivariate skew-normal, and multivariate Laplace distributions of the true generating process, respectively, with $n_{i}=500$ and $g=2$.
S5 Application: Microarray Data of Breast Cancer Patients
The dataset studied is taken from transcript profiles of primary breast tumors to investigate the p53 status. It is based on 251 samples predominantly containing tumor tissue that underwent sequencing to identify the p53 mutation status through RNA extraction and subsequent microarray analysis. The RNA was extracted from each sample by using Affymetrix U133 oligonucleotide microarrays, targeting over 30,000 genes (miller2005expression). Here we focus on a subset of genes corresponding to a signature that is considered important in breast cancer from the research of bonavita2020antagonistic, which stores the appropriately normalized expression values of $p=22$ gene products. The samples were divided into three groups based on tumor grade. After excluding two samples due to the missing grade variable, we obtain $g=3$ groups with $n_{1}=67$, $n_{2}=128$ and $n_{3}=54$ samples for grade I, grade II and grade III, respectively. Within a high dimensional framework, we have $p/\min(n_{i})=0.407$. Homoscedastic one-way MANOVA leads to the following $p$-values: $1.565\times 10^{-7}$ for the directional test, $1.966\times 10^{-12}$ for the central limit theorem test, $2.627\times 10^{-8}$ for the likelihood ratio test, $1.915\times 10^{-7}$ for the Bartlett corrected test, and $1.828\times 10^{-7}$ and $1.877\times 10^{-7}$ for the two Skovgaard’s modifications. All methods clearly indicate to reject the null hypothesis of equal breast tumors gene expression signature in case of homoscedastic groups. As already seen, the likelihood ratio test results in the smallest $p$-value, whereas its Skovgaard’s modified versions and the Bartlett corrected test are closer to the exact directional $p$-value, consistent with our simulations. To check the reliability of the various methods in the homoscedastic case, another simulation study on the gene expression data can be conducted. The empirical size based on $10,000$ replications for the directional, central limit theorem test, likelihood ratio, Bartlett corrected and two Skovgaard’s modifications tests are 0.049, 0.060, 0.087, 0.049 0.048 and 0.047, respectively. Those of central limit theorem test and likelihood ratio test are a little inflated. The remaining approaches, including the directional test, appear to be accurate.
Assuming heteroscedastic groups, for hypothesis (9) the directional $p$-value is $2.861\times 10^{-5}$, the $p$-value of the likelihood ratio test is $1.928\times 10^{-7}$, and the $p$-values of the two Skovgaard’s modifications are $1.772\times 10^{-5}$ and $2.079\times 10^{-5}$. The same conclusion is hence drawn by all methods, but once again the $p$-value obtained from the likelihood ratio test is relatively smaller than the others. Such discrepancies are particularly relevant in biostatistical problems, and more generally in medical research, where it is important to look at the specific size of $p$-values to get an idea about the clinical implications and relevance of the findings for the population of patients of interest. Even under heterescedasticity a simulation study can be performed. The empirical sizes of the directional test, the likelihood ratio test and the two Skovgaard’s modifications based on $10,000$ replications are 0.051, 0.181, 0.057 and 0.054, respectively. The directional test performs very well, as expected, while the likelihood ratio test seems still not reliable.